# Deep polytopic autoencoders for low-dimensional linear parameter-varying approximations and nonlinear feedback design

### v1

@@ Abstract
Polytopic autoencoders provide low-dimensional parametrizations of states in a polytope. For nonlinear PDEs, this is readily applied to low-dimensional linear parameter-varying (LPV) approximations as they have been exploited for efficient nonlinear controller design via series expansions of the solution to the state-dependent Riccati equation. In this work, we develop a polytopic autoencoder for control applications and show how it outperforms standard linear approaches in view of LPV approximations of nonlinear systems and how the particular architecture enables higher order series expansions at little extra computational effort. We illustrate the properties and potentials of this approach to computational nonlinear controller design for large-scale systems with a thorough numerical study.
@@ 1 Introduction
The design of controllers for nonlinear high-dimensional systems is a challenging task and there exists no generally applicable and computationally feasible approach. Here, we consider input-affine systems of the form
$\dot{v}(t)=f(v(t))+Bu(t),\quad y(t)=Cv(t),$ (1)
where for time $t>0$, $v(t)\in\mathbb{R}^{n}$ denotes the state, $u(t)\in\mathbb{R}^{m}$ and $y(t)\in\mathbb{R}^{\ell}$ denote the system’s input and output, $f\colon\mathbb{R}^{n}\to\mathbb{R}^{n}$ is a possibly nonlinear function, and $B$ and $C$ are linear input and output operators.
For the control of systems like eq. 1, methods like backstepping [17], feedback linearization [19, Chap. 5.3] and sliding mode control [12] require structural assumptions whereas general approaches based on the Hamilton-Jacobi-Bellman (HJB) equations are only feasible for very moderate system sizes; see, e.g., [8]. The commonly employed state-dependent Riccati equation (SDRE) approximation to the HJB (see, e.g., [10] for a survey) lifts a significant portion of the computational complexity. Nonetheless, this approach became available for large-scale systems only in recently due to further simplifications; see [6] for a linear update scheme and [1] for approximations through series expansions.
In this work, we expand on the series expansion of the SDRE combined with the approximation of the nonlinear system through a low-dimensional linear parameter-varying (LPV) system. This allows us to provide a generic algorithm to nonlinear controller design; see our previous works [16, 11] and related discussions in [1, 2]. In particular, the results in [16] indicate a measurable improvement of the nonlinear feedback design over a related, well-tuned and robust LQR controller but left space for further performance improvements, for example, through
  1. 1.
    higher order expansions of the nonlinear feedback or
  2. 2.
@@
Both topics are linked to the question if a nonlinear parametrization can provide better low-dimensional approximations with fewer degrees of freedom and thus mitigating the natural limit, which is commonly expressed through the Kolmogorov $n$-width of any linear model order reduction scheme – like proper orthogonal decomposition (POD) which is a linearly optimal projection scheme; [13]. There have been many efforts and positive results concerning the use of nonlinear model order reduction approaches using, e.g., polynomial manifold maps [9], learning techniques [13], or the selection of local bases [3].
To consider higher order expansions of a nonlinear feedback law, a very low-dimensional parametrization is crucial. Among others, as we will lay out, the numbers of coefficient that need to be precomputed through large-scale matrix equations equals $\sum_{k=0}^{p}\binom{r+k-1}{k}$, where $r$ is the dimension of the parametrization and $p$ is the order of the expansion. For example, for $r=10$ and a first-order expansion, the solve of $1+10=11$ matrix equations was required, cp. [16], whereas a second or third-order expansion would require $66$ or $286$ solves. However, for a parametrization of size $r=5$, only $6$, $21$, or $56$ matrix equation solves are needed for expansions of first, second and third order, respectively.
In this work, we rely on our recent development of polytopic autoencoders [14] that combines learning approaches with the selection of local bases in terms of polytope vertices. The concept of reconstructing the states in a polytope provides certain advantages over the reconstruction in a linear space. Generally, the observed states of a dynamical system are described in a bounded set rather than in an unbounded space. With the polytopic approach, the reconstruction is confined to a region where the states have been observed. This confinement stands against the freedom to explore the full state space, for controller design, however, it can be beneficial if tailored towards target regimes and since the resulting series expressions will be bounded by design.
@@
The underlying ideas and major findings of this paper can be summarized as follows: We recall in Section 2 that a general nonlinear control system can be approximated by a low-dimensional LPV system, which can be exploited for controller design in particular if the parameter dependency itself is linear. In view of providing a performant but maximally low-dimensional parametrization, we resort to clustered polytopic autoencoders that encode the states of a dynamical system in a polytope and decode them with locally adapted bases; see Section 3. An interesting result in view of controller design is that this nonlinear decoding has a Taylor expansion with vanishing second-order terms (Lemma 1). For the actual controller design in Section 4, we consider series expansions of state-dependent Riccati equation as it has been proposed and exercised before. Here, we extend the formulas to consider nonlinear parameter dependencies and higher order terms in the expansion. Finally, we report numerical experiments that illustrate the performance improvement of the polytopic autoencoder over standard approximation schemes, that explore the different levels of approximation to the underlying nonlinear SDRE feedback design, and that investigate the gain in controller performance for a commonly considered flow control problem; see Section 6.
2 Low-dimensional LPV approximations
@@
$\dot{v}(t)=A\bigl{(}\rho(v(t))\bigr{)}\,v(t)+Bu(t),\quad y(t)=Cv(t).$ (2)
Such systems can be derived from a state-dependent coefficient representation such as $f(v)=\widetilde{A}(v)v$ and a parametrization of the state $v(t)\approx\tilde{v}(t)=\nu(\mu(v(t))$, where $\mu$ is an encoder and $\nu$ is a decoder. In this case, we set $\rho=\mu(v)$ and $A(\rho):=\widetilde{A}(\nu(\rho))$. Note that the parameter depends on the state, hence the term quasi-LPV.
@@
In the example application we consider later, namely flows, we will work with coefficients $\widetilde{A}$ that are linear in the variable so that linear structures in the decoding $\rho\to\tilde{v}$ will lead to linear parameter dependencies as they have been proven beneficial for controller design; see, e.g., [4].
3 Polytopic parametrizations in clusters for controller design
@@
As opposed to linear model order reduction approaches, where the state is parametrized in a linear subspace, in polytopic autoencoders (PAEs), the state is represented through coordinates with respect to vertices of an $R$-dimensional polytope. As a polytope is bounded, without loss of generality, the vertices $\tilde{v}_{k}$ can be scaled and shifted such that the coordinates $\theta(t)\in\mathbb{R}^{R}$ that represent the reconstruction $\tilde{v}(t)$ in the polytope will be positive and sum up to one, i.e.,
$\tilde{v}(t)=\sum_{k=1}^{R}\theta_{k}(t)\tilde{v}_{k},\quad\text{with}\leavevmode\nobreak\ \theta(t)\geq 0,\quad\sum_{k=1}^{R}\theta_{k}(t)=1.$
@@
This design is not capable of exactly parametrizing a target value $v^{*}=0$ unless the support vectors $\tilde{v}_{k}$ come with a specific linear dependency. To explicitly enable the reconstruction of the zero target value and to ensure that the series expansion of the SDRE is about $\theta(\tilde{v})=0$, we include $\tilde{v}_{0}=0$ in the reconstruction basis of the training, where we impose the positivity and summation constraint on the $\theta$ variable extended by $\theta_{0}$, i.e.,
$\theta_{0}(t)+\sum_{k=1}^{R}\theta_{k}(t)=1.$ (3)
@@
This way, the tuple $(1,\theta(t))=(1,0)\in\mathbb{R}^{R+1}$ will be a feasible value in the extended parameter space that represents the target value.
Following previous results that showed how clustering, i.e., how the choice of local bases can improve the reconstruction, we combine the polytopic parametrization with smooth clustering networks as we have developed it in [14]. In short, we will use coordinates $\theta=c(\rho)\otimes\rho$, where $\rho\in\mathbb{R}^{r}$ denotes the actual latent variable that encodes $v$ and where $c$ is a smooth clustering network that blends between $q$ sets of local bases for the reconstruction.
Taking into consideration a zero target value, we propose a polytopic autoencoder that consists of three components:
  1. 1.
@@
    $(\rho_{0},\rho)=\mu(v)=:\operatorname{softmax}(\operatorname{F_{e}}(v)+e_{1}),$
    with a scalar function $\rho_{0}$ and $\rho(t)\in\mathbb{R}^{r}$, where the function $\operatorname{F_{e}}\colon\mathbb{R}^{n}\to\mathbb{R}^{r+1}$ is a trainable nonlinear convolutional neural network with $\operatorname{F_{e}}(0)=0$ and where $e_{1}\in\mathbb{R}^{r+1}$ is the first unit vector. Since $\operatorname{F_{e}}(0)=0$ and with the softmax function
    $\operatorname{softmax}\colon x\to\left(\tfrac{x_{i}\tanh(ax_{i})}{\sum_{j=1}^{r+1}x_{j}\tanh(ax_{j})}\right)_{i}^{r+1},$
    for a parameter $a>0$ (that we will set to $a=10$ in the experiments), we ensure that the output $(\rho_{0}(t),\rho(t))$ satisfies
    $(\rho_{0}(t),\rho_{k}(t))\geq 0\leavevmode\nobreak\ \leavevmode\nobreak\ \text{and}\leavevmode\nobreak\ \leavevmode\nobreak\ \rho_{0}(t)+\sum_{k=1}^{r}\rho_{k}(t)=1;$
    in particular $\mu(0)=(1,0)$.
  2. 2.
    A smooth clustering network $c\colon\mathbb{R}^{r}\to\mathbb{R}^{q}$ so that
@@
    where $\operatorname{F_{c}}$ is a trainable neural network with $\operatorname{F_{c}}(0)=0$ and where $\hat{e}_{1}$ is the first unit vector in $\mathbb{R}^{q}$. By design, for the output $c(\rho)$, we have that
@@
    hold, so that $(\theta_{0},\theta):=(\rho_{0},c(\rho)\otimes\rho)$ satisfies the positivity and summation conditions eq. 3.
  3. 3.
    A trainable linear map $W\in\mathbb{R}^{n\times qr}$ that realizes the reconstruction as
    $\tilde{v}=W\big{(}c(\rho)\otimes\rho\big{)}.$ (5)
We lay out the design and motivation in the following remarks:
  • The property that $\operatorname{F_{e}}(0)=0$ and $\operatorname{F_{c}}(0)=0$ is achieved by using (convolutional) nonlinear layers without bias terms and with activation functions that cross the origin.
@@
  • The design of the clustering network ensures $c(0)=e_{1}$ so that values of $\rho$ close to zero are associated with the first cluster and that $c(\mu(0))\otimes\mu(0)=(1,0)\in\mathbb{R}^{qr}$ represents the target vector $v^{*}=0$ exactly.
  • With using the Kronecker product $c(\rho)\otimes\rho$, the $k$-th $r$ columns of $W$ denote the supporting basis for reconstruction in the $k$-th cluster; see [14].
@@
With $\operatorname{softmax}\colon x\to\left(\tfrac{x_{i}\tanh(ax_{i})}{\sum_{j=1}^{q}x_{j}\tanh(ax_{j})}\right)_{i}^{q}$, it holds that the Jacobian $c^{\prime}$ of the clustering network $c$ as defined in eq. 4 is zero at $\rho=0$. In particular, the second-order Taylor expansion of $\rho\to A(\rho)$ coincides with the first-order expansion and reads
@@
By direct computations we find $c^{\prime}(\operatorname{F_{c}}(\rho)+\hat{e}_{1})\big{\rvert}_{\rho=0}=c^{\prime}(\hat{e}_{1})=0\in\mathbb{R}^{q\times r}$. Since the coefficient map $c(\rho)\otimes\rho\to\widetilde{A}(\tilde{v}(\rho))$ is linear, we can reduce the argument to the nonlinear part $\rho\to c(\rho)\otimes\rho$. With $\rho=0$, from
@@
$c(0+h)\otimes h=\left(c(0)+c^{\prime}(0)h+o(\lvert h\rvert^{2})\right)\otimes h$
and as
@@
where $c^{\prime}(0)\in\mathbb{R}^{q\times r}$ is the Jacobian of $c$ at $\rho=0$. With $c^{\prime}(0)=0$, the second-order terms vanish so that with $c(0)=\hat{e}_{1}$, the first (and second-order) approximation to the polytopic reconstruction reads
$\tilde{v}(\rho)\approx W\left(\hat{e}_{1}\otimes\rho\right)=\sum_{j=1}^{r}\rho_{j}w_{j}.$ (7)
Therefore, formula eq. 6 follows from the linearity of $\widetilde{A}$. ∎

### v2

# 2025-01-23

@@ Abstract
Polytopic autoencoders provide low-dimensional parametrizations of states in a polytope. For nonlinear PDEs, this is readily applied to low-dimensional linear parameter-varying (LPV) approximations as they have been exploited for efficient nonlinear controller design via series expansions of the solution to the state-dependent Riccati equation. In this work, we develop a polytopic autoencoder for control applications and show how it improves on standard linear approaches in view of LPV approximations of nonlinear systems. We discuss how the particular architecture enables exact representation of target states and higher order series expansions of the nonlinear feedback law at little extra computational effort in the online phase and how the linear though high-dimensional and nonstandard Lyapunov equations are efficiently computed during the offline phase. In a numerical study, we illustrate the procedure and how this approach can reliably outperform the standard linear-quadratic regulator design.
keywords:
Nonlinear systems, nonlinear feedback control, parameter-varying approximations, Riccati equations, Lyapunov equations
@@ 1 Introduction
The design of feedback controllers for nonlinear high-dimensional systems is a challenging task and there exists no generally applicable and computationally feasible approach. Here, we consider input-affine systems of the form
$\dot{x}(t)=f(x(t))+Bu(t),\quad y(t)=Cx(t),$ (1)
where for time $t\geq 0$, $x(t)\in\mathbb{R}^{n}$ denotes the state, $u(t)\in\mathbb{R}^{m}$ and $y(t)\in\mathbb{R}^{\ell}$ denote the system’s input and output, $f\colon\mathbb{R}^{n}\to\mathbb{R}^{n}$ is a possibly nonlinear function, and $B\in\mathbb{R}^{n\times m}$ and $C\in\mathbb{R}^{p\times n}$ are linear input and output operators represented as matrices. The general idea of feedback controller design is to identify a mapping $\mathcal{K}\colon x(t)\to u(t)$ so that with the choice of $u=\mathcal{K}(x)$, the system’s state $x$ and output $y$ satisfy desired criteria. A typical application is feedback stabilization that seeks to achieve $x(t)\to 0$ as $t\to\infty$ or the more general task of set-point control, namely $x(t)\to x^{\ast}$ for some state $x^{\ast}\in\mathbb{R}^{n}$ that fulfills $f(x^{\ast})=0$.
For the controller design for systems of the form eq. 1, methods like backstepping [34], feedback linearization [41, Chap. 5.3], and sliding mode control [22] require structural assumptions, whereas general approaches based on the Hamilton-Jacobi-Bellman (HJB) equations are only feasible for very moderate system sizes or in combination with model reduction; see, e.g., [16] for polynomial expansions, [23] for tensorized approximations, or [3] for an approach that addresses general PDEs based on first reducing the state-space dimension through, e.g., POD. Other efforts try to approximate the solution of the HJB equations by deep neural networks; see [20]. The commonly employed state-dependent Riccati equation (SDRE) approximation to the HJB equations (see, e.g., [18] for a survey) lifts a significant portion of the computational complexity. Nonetheless, this approach became available for large-scale systems only recently due to further simplifications; see [13] for a linear update scheme and [1] for approximations through series expansions for systems of a certain parametric structure. These structural assumptions have been lifted in our previous work (see [30]) so that, with the help of model reduction, now a large class of nonlinear control systems can be treated with series expansions of the SDRE.
The underlying idea is the approximation of the nonlinear system eq. 1 through a very low-dimensional linear parameter-varying (LPV) system so that the series expansion of any nonlinear feedback law can be formulated with respect to a few parameters rather than the full state; see our previous works [30, 21] and related discussions in [1, 2]. In particular, the results in [30] indicate a measurable improvement of the nonlinear SDRE feedback over a related, well-tuned and robust linear-quadratic regulator (LQR) but left space for further performance improvements, for example, through
  • (i)
    higher order expansions of the nonlinear feedback, or
  • (ii)
@@
For higher order expansions of nonlinear feedback laws, a low-dimensional parametrization of the states is crucial since the number of coefficients grows drastically with the dimension of the parametrization and the order of the expansion so that both topics are linked to the single question if a nonlinear parametrization can provide sufficiently accurate low-dimensional approximations with only few degrees of freedom.
It is known that linear parametrizations like the (linearly) optimal approach of proper orthogonal decomposition (POD, see, e.g., [26]) have a natural limit on the approximation accuracy, which is commonly expressed through the Kolmogorov $n$-width; see [37]. Thus, there have been many efforts and positive results concerning the use of nonlinear model order reduction approaches including, for example, polynomial manifold maps [9, 17] or general nonlinear autoencoders [26, 40, 38]. An alternative for accurate low-dimensional parametrizations is found in the use of local bases [4]; see the recent works that consider local POD bases state reconstruction on subsets of the given dataset [5, 6, 33] or our previous work on local bases approximations via autoencoders [32, 31]. Basically, local bases are designed for subsets of states considered to have high similarities. However, local bases require a mechanism to identify the subsets and which basis is to be used. Standard approaches use clustering for this purpose, for example, $k$-means clustering, fuzzy $c$-means clustering, and deep clustering [8, 24, 25].
In this work, we call on autoencoders with smooth basis selection and present a purely computational approach to design reduced-order parametrizations of nonlinear equations and provide algorithms and basic results for applications in nonlinear feedback control. Therefore, we extend our recently developed framework of polytopic autoencoders [32] that combines an autoencoder with the selection of local bases in terms of polytope vertices towards the use in set-point control. In particular, we directly impose the set-point condition $f(x^{\ast})=0$ onto the autoencoder so that we obtain a nonlinear encoding of very small dimension with arbitrary accuracy in the relevant part of the state space. To this end, we derive explicit expressions for the Jacobian of the decoder, which we then can exploit in theory and practice for analysis and synthesis of higher order expansions of the SDRE feedback. The precomputation of the expansion coefficients requires the solution of large-scale matrix-valued algebraic equations of a nonstandard form. Here, we call and expand on recent advances in the solution of indefinite matrix Lyapunov and Riccati equations; see [35, 39]. We note that the derived representation through a polytopic LPV system is suited for controller design on the base of linear matrix inequalities (LMIs); see, e.g., [7]. However, since the reduction only considers the structure but not the high-dimensional state space, the systems to solve are far too large for state-of-the-art LMI solvers. We exemplify the procedure and potentials of this general-purpose approach to nonlinear feedback design with a numerical example based on the two-dimensional flow past a cylinder.
@@
The overall structure of this paper can be summarized as follows: We recall in Section 2 that a general nonlinear control system can be approximated by a low-dimensional LPV system, which can be exploited for controller design in particular if the parameter dependency itself is linear. In view of providing a performant but maximally low-dimensional parametrization, we resort to clustered polytopic autoencoders that encode the states of a dynamical system in a polytope and decode them with locally adapted bases; see Section 3. An interesting result in view of controller design is that this nonlinear decoding has a Taylor expansion with vanishing second-order terms (Lemma 1). For the actual controller design in Section 4, we consider series expansions of state-dependent Riccati equation. Here, we extend the formulas to consider nonlinear parameter dependencies and higher order terms in the expansion. We discuss and showcase, how the relevant large-scale matrix equations can be solved numerically. Finally, we report numerical experiments that illustrate the performance improvement of the polytopic autoencoder over standard approximation schemes that explore the different levels of approximation to the underlying nonlinear SDRE feedback design, and that investigate the gain in controller performance for a commonly considered flow control problem; see Section 6. As a result, we obtain a model-based nonlinear controller of reduced complexity, with the reduction tailored to snapshot data. The paper is concluded in Section 7.
2 Low-dimensional linear parameter-varying approximations
@@
$\dot{x}(t)=A\bigl{(}\rho(x(t))\bigr{)}x(t)+Bu(t),\quad y(t)=Cx(t).$ (2)
Such systems can be derived from an (exact) state-dependent coefficient representation such as $f(x)=\widetilde{A}(x)x$ and a parametrization of the state $x(t)\approx\tilde{x}(t)=\nu(\mu(x(t))$, where $\mu$ is an encoder and $\nu$ is a decoder; see [30]. In this case, we set $\rho=\mu(x)$ and $A(\rho):=\widetilde{A}(\nu(\rho))$. Note that the parameter $\rho$ depends on the state, hence the term quasi-LPV.
@@
In the example application of fluid flow models, the coefficient function $\widetilde{A}$ is readily available; cf. [28]. Here, the function $\widetilde{A}$ is linear so that a linear decoding will lead to linear parameter dependencies as they have been proven beneficial for controller design; see, e.g., [7].
In what follows, we refer by $\rho_{i}$ to the $i$-th component of a vector-valued quantity $\rho$. To reduce the notation, we omit dependencies and sometimes write, e.g., $\rho$ or $\rho(t)$ instead of $\rho(x(t))$. Furthermore, we will denote the Kronecker product by $\otimes$.
3 Polytopic parametrizations for control
@@
In polytopic autoencoders (PAEs), the state is represented through generalized barycentric coordinates with respect to $R\in\mathbb{N}$ vertices $\{\tilde{x}_{k}\}_{k=1}^{R}\subset\mathbb{R}^{n}$ of a polytope; see, e.g., our previous work [32] for a more detailed introduction. As a polytope is bounded, one can always assume without loss of generality that the coordinates $\rho(t)\in\mathbb{R}^{R}$ in the reconstruction will be positive and sum up to one, i.e., for the $\tilde{x}(t)$ in the polytope it holds that
$\tilde{x}(t)=\sum_{k=1}^{R}\rho_{k}(t)\tilde{x}_{k},\quad\text{with }\rho(t)\geq 0\text{ and }\sum_{k=1}^{R}\rho_{k}(t)=1.$ (3)
An immediate advantage of the polytopic parametrization for the intended controller design through series expansions results from the boundedness of the reconstruction, which keeps the controller gains bounded so that we can expect good approximations in series expansions while possible scaling issues will be detected when precomputing the expansion coefficients.
@@
For the application in controller design, the consistency at the target state $x^{\ast}$ is important. We note that a design like eq. 3 is not capable of exactly parametrizing a target value $x^{\ast}=0$ unless the vertices $\tilde{x}_{k}$ come with a specific linear dependency. Moreover, for the intended series expansions around the target state, accuracy is guaranteed if $\|\rho(t)\|$ is small, which, however, contradicts the summation condition in (3). To address these two requirements, we include $\tilde{x}_{0}=0$ in the training of the reconstruction basis, where we impose the positivity and summation constraint on the $\rho$ variable extended by $\rho_{0}$, i.e.,
$\rho_{0}(t)+\sum_{k=1}^{R}\rho_{k}(t)=1.$ (4)
@@
This way, the tuple $(1,\rho(t))=(1,0)\in\mathbb{R}^{R+1}$ will be a feasible value in the extended parameter space that exactly represents the setpoint $x^{\ast}=0\in\mathbb{R}^{n}$. Moreover, smoothness in the parameter space, will ensure that $\rho$ is small around the origin.
As for the use of local bases, we assume that $q$ different polytopes with vertices
$\{\tilde{x}_{k}^{(j)}\}_{k=1,\dotsc,R;\,j=1,\dotsc,q}\subset\mathbb{R}^{n}$
are to be used as local bases for the reconstruction. Then, a direct implementation would compute both the local coordinates $\rho$ and the index $j$ of the relevant polytope and reconstruct according to the formula
$\tilde{x}=\begin{bmatrix}\tilde{x}_{1}^{(1)}&\ldots&\tilde{x}_{R}^{(1)}&\tilde{x}_{1}^{(2)}&\ldots&\tilde{x}_{R}^{(2)}&\ldots&\tilde{x}_{1}^{(q)}&\ldots&\tilde{x}_{R}^{(q)}\end{bmatrix}(\alpha\otimes\rho)$ (5)
where, in this case, $\alpha=e_{j}$ is the $j$-th canonical basis vector. Instead of this hard selection of the $j$-th polytope, we rather use a smooth clustering network (e.g., [43, 25]) that blends between the polytopes by choosing $\alpha=c(\rho)\in\mathbb{R}^{q}$, with $c(\rho)\geq 0$ and $\sum_{\ell=1}^{q}c(\rho)_{\ell}=1$. We note that the summation condition on $c(\rho)$ allows to consider $c(\rho)\otimes\rho$ as generalized barycentric coordinates in the polytope spanned by all vertices as in eq. 5; see [32, Lem 3.1].
Under the preceding considerations, we propose a polytopic autoencoder designed for set-point control towards $x^{\ast}=0$ that consists of the following three components:
  1. (i)
@@
    $(\rho_{0},\rho)=\mu(x)=:\operatorname{softmax}(\operatorname{F_{e}}(x)+e_{1}),$
    with a scalar function $\rho_{0}$ and $\rho(t)\in\mathbb{R}^{r}$, where $e_{1}\in\mathbb{R}^{r+1}$ is the first unit vector and $\operatorname{F_{e}}\colon\mathbb{R}^{n}\to\mathbb{R}^{r+1}$ is a nonlinear convolutional neural network described via
    $h_{1}$ $=\mathsf{T}x,$
    $h_{2}$ $=\mathsf{ELU}(\mathsf{Conv}_{1}(h_{1})),$
[…]
    $h_{L}$ $=\mathsf{ELU}(\mathsf{Conv}_{L-1}(h_{L-1})),$
    $\operatorname{F_{e}}(x)$ $=\mathsf{FC}(h_{L}).$
    Here, $\mathsf{T}$ denotes the interpolation of spatially distributed data onto a pixel grid as it is needed for the convolutions, $\mathsf{ELU}$ is the exponential linear unit (ELU) activation function [19], $\mathsf{FC}$ is a fully connected layer, and $\mathsf{Conv}_{i}$ is a convolutional layer, $i\in\{1,2,\cdots,L-1\}$ for some depth parameter $L\in\mathbb{N}$. Each layer is designed without bias terms to ensure $\operatorname{F_{e}}(0)=0$ so that with the softmax function
    $\operatorname{softmax}(\xi)_{i}=\tfrac{\xi_{i}\tanh(10\xi_{i})}{\sum_{j=1}^{r+1}\xi_{j}\tanh(10\xi_{j})}$
    for $i=1,\ldots,r+1$ and $\xi=(\xi_{1},\ldots,\xi_{r+1})\in\mathbb{R}^{r+1}$, we achieve, by design, that the output $(\rho_{0}(t),\rho(t))$ satisfies
    $(\rho_{0}(t),\rho_{k}(t))\geq 0\quad\text{and}\quad\rho_{0}(t)+\sum_{k=1}^{r}\rho_{k}(t)=1;$
    and in particular $\mu(0)=(1,0)$.
  2. (ii)
    A smooth clustering network $c\colon\mathbb{R}^{r}\to\mathbb{R}^{q}$, with
@@
    where $\hat{e}_{1}$ is the first unit vector in $\mathbb{R}^{q}$ and
    $\operatorname{F_{c}}(\rho)=\mathsf{ELU}(\mathsf{FC}(\rho)),$
    with no bias, i.e., $\operatorname{F_{c}}(0)=0$. By design, for the output $c(\rho)$, we have that
@@
    hold, so that $(\rho_{0},c(\rho)\otimes\rho)$ satisfies the positivity and summation conditions eq. 4.
  3. (iii)
    A trainable linear map $W\in\mathbb{R}^{n\times qr}$ that realizes the reconstruction (cf. eq. 5) as
    $\tilde{x}=W\big{(}c(\rho)\otimes\rho\big{)}.$ (7)
A few comments on the design of the autoencoder and motivation:
  • Generally, the property that $\operatorname{F_{e}}(0)=0$ and $\operatorname{F_{c}}(0)=0$ can be achieved by using standard (convolutional) linear layers without bias terms and with any activation function that crosses the origin.
@@
  • The smooth clustering network, as opposed to a standard clustering, also ensures differentiability of the network as it will become important for the series expansion.
  • The design of the clustering network ensures that $c(0)=e_{1}$ so that values of $\rho$ close to zero are associated with the first cluster and that $c(\mu(0))\otimes\mu(0)=(1,0)\in\mathbb{R}^{qr}$ represents the target vector $x^{\ast}=0$ exactly.
With the following lemma, we summarize the resulting properties when employing the autoencoder in series expansions of the coefficient function.
@@
Given the smooth clustering network $c$ as defined in eq. 6 and the corresponding LPV coefficient function
$\rho\to A(\rho)=\widetilde{A}\Bigr{(}W\,\bigl{(}c(\rho)\otimes\rho)\Bigr{)}$ (8)
defined through the linear map $\widetilde{A}\colon\mathbb{R}^{n}\to\mathbb{R}^{n\times n}$ and the decoder eq. 7. Then, it holds that
  • (a)
    the coefficient function eq. 8 is differentiable,
  • (b)
    the Jacobian of the clustering network $c$ is zero at $\rho=0$, i.e.,
    $c^{\prime}(\rho)\bigl{|}_{\rho=0}=0\in\mathbb{R}^{q,r},$
    and
  • (c)
    the second-order Taylor expansion of eq. 8 coincides with the first-order expansion and reads
@@
As for the result (a), we note that, by construction, $c$ is a concatenation of arbitrarily often differentiable functions. As for the part (b), by direct computations, we can confirm that $c^{\prime}(\operatorname{F_{c}}(\rho)+\hat{e}_{1})\big{\rvert}_{\rho=0}=c^{\prime}(\hat{e}_{1})=0\in\mathbb{R}^{q\times r}$.
Since the coefficient map
$\rho\to\widetilde{A}\Bigr{(}W\,\bigl{(}c(\rho)\otimes\rho)\Bigr{)}$
is linear in components that concern $W$ and $\widetilde{A}$ (cf. Equation 7 and Remark 1), we can reduce the argument to the nonlinear part $\rho\to c(\rho)\otimes\rho$. With $\rho=0$, from
@@
$c(0+h)\otimes h=\left(c(0)+c^{\prime}(0)h+\mathcal{R}(h)\right)\otimes h$
with the remainder term $\mathcal{R}(h)\in o(|h|)$, meaning that $\mathcal{R}$ approaches 00 faster than $h$ itself. Thus, the second-order expansion of $\rho\to c(\rho)\otimes\rho$ around the origin 00 reads
@@
where $c^{\prime}(0)\in\mathbb{R}^{q\times r}$ is the Jacobian of $c$ at $\rho=0$. With $c^{\prime}(0)=0$, the second-order terms vanish so that with $c(0)=\hat{e}_{1}$, the second-order expansions reduces to the first order approximation of the polytopic reconstruction that reads
$\tilde{x}(\rho)\approx W\left(c(0)\otimes\rho\right)=W\left(\hat{e}_{1}\otimes\rho\right)=W\begin{bmatrix}\rho\\ 0\\ \vdots\\ 0\end{bmatrix}=\sum_{j=1}^{r}\rho_{j}w_{j}.$
Finally, formula eq. 9 follows from the linearity of $\widetilde{A}$ and $W$. ∎
Remark 2.
We note the following consistency property in the smooth clustering. By construction, the first $r$ columns of $W$ describe the local basis for the reconstruction in the cluster associated with $\rho=0$. Accordingly, the linearization around the origin only considers the basis that is associated with the origin.